OpenStudy (gabyramirezzz):

Which function has the range (-infinity, -2]U[0,infinity)

OpenStudy (gabyramirezzz):

a) y= cos(x+1) b) y= sec(x) +1 c) y= csc(x) -1 d) y= cot(2x)-1

geerky42 (geerky42):

Can you tell what range is for a?

OpenStudy (gabyramirezzz):

all real numbers except multiples of pi?

OpenStudy (gabyramirezzz):

i think its either b or c im just not sure which

geerky42 (geerky42):

None, Do you know what graph of \(\cos(x+1)\) looks like? https://www.desmos.com/calculator/vhsoioyx9j Range is \([-1,1]\), right?

OpenStudy (gabyramirezzz):

right

geerky42 (geerky42):

What about b? do you know what is range of sec(x)? Now what's range of sec(x)+1?

OpenStudy (gabyramirezzz):

(-infinity, -1]U[1, infinity)

geerky42 (geerky42):

Yeah, that's range of sec(x) sec(x)+1 is same except that it is moved 1 unit up. So new range would be (-infinity, 0]U[2, infinity)

geerky42 (geerky42):

So now you can tell that answer is C.

OpenStudy (gabyramirezzz):

ohhh

OpenStudy (gabyramirezzz):

thank you!!

geerky42 (geerky42):

Great, glad I helped!

OpenStudy (anonymous):

no its b i know

geerky42 (geerky42):

1) No it's not b. See here: https://www.desmos.com/calculator/pb951bcx8d 2) This is 5 months old question. What are you doing here anyway?

OpenStudy (anonymous):

it is c